OpenStudy (anonymous):

Did i do this right? i had tan(27)=x/31 and for x I got 60.8409....

OpenStudy (anonymous):

\[tan(27)=\frac{x}{31}\] \[x=31tan(27)=15.8\] rounded

OpenStudy (anonymous):

i am assuming that 27 means 27 degrees

OpenStudy (anonymous):

oh okay but what about when the x is on bottom like cos(34degrees)=50/x?

OpenStudy (anonymous):

then solve like any ratio: \[cos(\theta)=\frac{50}{x}\] \[x=\frac{50}{cos(\theta)}\]

OpenStudy (anonymous):

in this case you get \[x=\frac{50}{cos(34)}=60.31\]

OpenStudy (anonymous):

ohh okay thanks!!<3